OpenStudy (anonymous):

I really need help with this, i dont know how to plug the 3 into the inverse. Medal for whoever helps :) Given f(x)= x - 5 over 3, solve for f^-1(3)

OpenStudy (misty1212):

solve \[\frac{x-5}{3}=3\] in two steps

OpenStudy (misty1212):

multiply by 3, add 5

OpenStudy (anonymous):

what do i multiply by 3?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

both sides dear you are just trying to clear the fraction

OpenStudy (misty1212):

\[\frac{x-5}{3}=3\\ x-5=9\]

OpenStudy (anonymous):

that cant be right though, my choices are: -1 14 18 and 21